OpenStudy (anonymous):

log abc^0.5+log abc =?

OpenStudy (anonymous):

options are: 5,9,2,1

OpenStudy (xapproachesinfinity):

let me see if this what you wrote before we begin \[\log abc^{\frac{1}{2}}+\log abc\]

OpenStudy (xapproachesinfinity):

is that right?

OpenStudy (xapproachesinfinity):

is it (abc)^0.5 or just ab(c^0.5) huge difference

OpenStudy (xapproachesinfinity):

so which is it?

OpenStudy (xapproachesinfinity):

hello!! you there?

OpenStudy (anonymous):

yes the first one is ryt

OpenStudy (anonymous):

will be 1/2log abc +log abc

OpenStudy (xapproachesinfinity):

so it is (abc)^1/2 yes

OpenStudy (anonymous):

yes

OpenStudy (xapproachesinfinity):

okay then \(\log (abc)^{frac{1}{2}}+\log abc= \frac{1}{2} \log abc+\log abc\) \(=\frac{3}{2}\log(abc)\)

OpenStudy (xapproachesinfinity):

that first line where frac12 supposed to be all to 1/2

OpenStudy (xapproachesinfinity):

so we get \(=\log(abc)^{\frac{3}{2}}\)

OpenStudy (anonymous):

yea , next step

OpenStudy (xapproachesinfinity):

that is all what is it you want to achieve?

OpenStudy (xapproachesinfinity):

a different way of doing that is \(\log (abc)^{\frac{1}{2}}+\log abc=\log \frac{(abc)^{\frac{1}{2}}}{abc}\)

OpenStudy (anonymous):

but the options of this ques given are a) 5 b) 9 c) 2 d) 1 which onw wud be the ans?

OpenStudy (xapproachesinfinity):

i meant \(=log[(abc)^0.5 (abc)]\) not division!!

OpenStudy (xapproachesinfinity):

\(=\log[(abc)^{0.5} (abc)]\)

OpenStudy (xapproachesinfinity):

if those are your options then you missing something in the question

OpenStudy (anonymous):

alright

OpenStudy (xapproachesinfinity):

can you post the picture of your question

OpenStudy (anonymous):

that must be the wrong ques. i pick this ques from mformaths.

OpenStudy (anonymous):

bdw thankew for your help. :)

OpenStudy (xapproachesinfinity):

that question cannot be any of those number no more simplification is required

OpenStudy (anonymous):

yep

OpenStudy (xapproachesinfinity):

all you can do is something like this \(=3/2[\log a+\log b+ \log c]\) which pretty much does not help for anything new

OpenStudy (anonymous):

ok i have another question cud u help me in that?

OpenStudy (xapproachesinfinity):

post is in a different post though

OpenStudy (anonymous):

okz